### v5

@@ Abstract
To find a suitable framework for nonlinear crystal optics(NCO), we have revisited linear crystal optics(LCO). At the methodological level, three widely used plane wave bases are compared in terms of eigenanalysis in reciprocal space and light field propagation in real space. Inspired by complex ray tracing, we expand M.V. Berry & M.R. Dennis’s 2003 uniform plane wave model to non-uniform Fourier crystal optics(FCO) and ultimately derive the explicit form of its 32 transition matrix, bridging the two major branches of crystal optics(CO) in reciprocal space, where either ray direction or spatial frequency serves as the input variable. Using this model, we create the material-matrix tetrahedral compass(M-M TC) to conduct a detailed analysis of how the four fundamental characteristics of materials (linear/circular birefringence/dichroism) influence the eigensystems of the vector electric field in two-dimensional(2D) spatial frequency domain and its distribution in three-dimensional(3D) space with a crystal-2f configuration. Along this journey, we have uncovered new territories in LCO in both real and reciprocal space, such as infinite singularities arranged in disk-, ring-, and crescent-like shapes, “L shorelines” resembling hearts, generalized haunting theorem, double conical refraction(DCR), and optical knots(OKs) it induces. We also present our model’s early applications in focal engineering and NCO. For focusing: dual-eigenmode (1) decomposition, explaining Raman spikes in conical refraction(CR); (2) their interference, leading to longitudinal multifoci; (3) backward propagation, for non-Zernike aberration correction. For NCO: (1) chiral second harmonic conical refraction(SHCR); (2) linear and nonlinear spin-orbit couples; (3) nonlinear optical orbital angular momentum(OAM) non-conservation — all three vector/scalar NCO phenomena, reproduced pixel by pixel. As the opening chapter in a trilogy, this work interweaves crystal optics, Fourier optics, and nonlinear optics, while integrating theoretical, computational, and experimental physics, advancing all six domains.
@@
Crystal optics, Singular optics, Wave optics, Fourier optics, Nonlinear optics, Quantum optics, Angular spectrum theory, Anisotropic diffraction, Plane wave model, Vector field, Maxwell equation, Partial differential equation, Constitutive relations, Boundary conditions, Generalized Snell’s law, Eigenanalysis, Characteristic equationContents
@@ 1 Introduction
Threefold purpose guides us: to establish crystal optics(CO) and Fourier optics(FO) as the physical and computational bases of nonlinear crystal optics(NCO) in the first part of a trilogy; to propose that three plane-wave bases form the mathematical core of CO spectral analysis, each offering trade-offs; to seek and knock the computational boundaries of linear crystal optics(LCO) and NCO via large-scale FO simulations in both three-dimensional(3D) reciprocal and two-dimensional(2D) spatial domains. With this, we proceed.
@@
Think about the ordinary(o) and exordinary(e) waves in nonlinear crystal optical(NCO) phase-matching types[6, 7, 8, 9, 10, 11, 12](e.g. type-I ooe in Fig. 5a): what are they? Eigenmodes governed by CO[13, 4, 14, 15, 16, 17, 18, 19]!! That is to say, the entire NCO process is constantly subject to the constrains imposed by CO[13, 5, 3, 4], which is itself inherently complex[20]. More phenomenologically, for all frequency-mixing phenomena, both the pumps/fundamental waves(FWs) and the newly generated frequencies(NFs) within the crystal, must be first decomposed into the material’s intrinsic eigenmodes, then diffract independently (and anisotropically) during subsequently linearly superposition and interference, while potentially enduring absorption or gain[21, 22].
@@
Given that the impact of LCO(= LO CO) on NCO has been historically either ignored or modeled incorrectly[22], we undertake a thorough reexamination of LCO in Supplementary Note 1. Through this inquiry, we conclude that:
@@
Only two options remain (see “Methods”): uniform CO basis that contradicts both FO and boundary conditions(BCs), but only requires solving bi-quadratic equations thanks to spherical coordinates(); rectangular() non-uniform CO basis that fits both physics(FO) and experiments(BCs), but demands batchly solving quartics with intractably long formulas(exceeds a single A4 page).
@@
In this Article, we propose a non-uniform linear Fourier crystal optical(LFCO) model, whose CO part integrates the advantages of both uniform[75, 76, 77, 78, 79, 80, 81, 1, 5, 20] and non-uniform[83, 15, 85, 14, 84, 86, 87, 111, 112] plane wave models , solving a pure bi-quadratic under FO and boundary constraints. This LFCO model extends the standard non-diagonal/non-positive-definite[113, 114, 115, 116, 117, 118, 119, 113, 120, 121, 122]/non-Hermitian[123, 124, 125, 126, 127, 128, 129, 130, 131, 132]/non-unitary[133, 134, 135, 136]/non-normal[107]/non-(mirror-[137])symmetric[5, 138, 139]/non-reciprocal[140, 141, 137] 22[142] transfer[98, 143]/transition[98, 110] matrix( evolution operator[144]) in CO to a 32 form (see Eqs. 5 and 6), enabling explicit fully vectorial[93, 91, 145, 146, 147, 148, 149, 150, 151] computation within arbitrary materials, and at the same time, allows the LO part to directly operate in the non-uniform[83, 15, 85, 14, 84, 86, 87, 92, 84] FO framework[91, 62, 90, 93, 89].
@@
To indirectly validate this LFCO model, and highlight its powerful role in advancing the analytic development of NCO, we offer a roadmap for expanding the LFCO model in Fig. 2c. Following its conceptual architectures — scalar(, semi-vector,) and (full-)vector NFCO — we first demonstrate this LFCO model’s applications: in full-vector NFCO, via chiral second-harmonic conical refraction(SHCR) in Figs. 1 and 2, and harmonic spin-orbit(S-O) angular momentum(AM) cascade in Fig. 3 (see also Figs. SC3-SC5); in scalar NFCO, through phase-matching-controlled orbital angular momentum(OAM) conversion in Fig. 4 (see also Figs. SB2-SB6); and in semi-vector NFCO, via full conical phase matching(FCPM) in Fig. 5 (see also Fig. SC2).
@@
To directly validate this LFCO model, as stated at the beginning of Section 2.4, the procedure must begin with verifying the correctness of the eigenvalue-eigenvector computations (Fig. S8), followed by confirming the accuracy of the field distributions in real space (Fig. 7, Fig. 8, Figs. S9-S11, and Figs. SA2-SA10).
@@
Accordingly, we introduce the material-matrix(M-M) tetrahedron compass(TC) in Fig. S8, where scanning parameters along three of its edges depict a theoretical panorama of the adiabatic evolution(AE) of electric field eigenmodes in the domain[152], competing among the three primary material properties, i.e. linear dichroism(LD), circular dichroism(CD), and optical activity(OA). During this process, we observe that CD, adhering to the haunting (C points) theorem[5] as well, can lead to heart-shaped L shorelines (as an upgrade to the L lines in CO[5]) and an infinite array of singularities in 2D domain, often arranged in patterns resembling disks, rings[101, 153], or crescents (instead of 8 finite EPs[5, 154, 154, 155, 22, 92, 156]). In Fig. S8, certain known phenomena reappeared[101, 153, 157], while the others, to the best of our knowledge, are original.
@@
Aided by crystal-2f system proposed from Fig. S9, we map out a second, more experimentally relevant[158, 159, 139, 81, 160] panorama in the 3D domain in Fig. 7, freezing the AE of the light field distribution between the three vertices of the M-M TC, corresponding to the material’s birefringence(Bi), LD, and OA.
@@
Still in real space, but with higher numerical aperture(N.A.), we offer a unified solution for the forward propagation and inverse design of focal fields in Fig. 8, where a dual-eigenmode decomposition is proposed to explain Raman spikes[161] in conical refraction(CR) arising solely from slow modes, multifoci induced by laser processing inside materials, and aberration correction without Zernike polynomials. Along the way, we have observed two surprising new phenomena: double conical refraction(DCR) and the resulting optical field knots.
@@
Owing to the intrinsic breadth, depth, and informational load, the main text cannot be fully decoupled from the Supplementary Notes, but we’ve maximized clarity to help readers — particularly in NCO — quickly grasp our LCO methods, results, and conclusions, which greatly benefit NCO.
This work unites mathematical elegance(see Supplementary Notes 2-5), physical intuition(see Supplementary Notes 7), and experimental accuracy(see Supplementary Note 8-9). Even so, it too suffers computational limits(see Supplementary Notes 5-6).
@@ 2 Results
2.1 By 2025, can/has anyone reproduce(d) chiral SHCR?
@@
The optically active( chiral[160]) SHCR[159, 165, 166, 162, 163] experiment in Fig. 1f[162] represents an NCO experiment that, in principle, cannot be well-reproduced without a mature LCO framework[163, 164, 167, 168, 166, 169, 170, 171, 162, 172, 173]. All previous NCO modeling attempts[170, 168, 168, 166, 167] have invariably failed to capture the experimental phenomena at a pixel-level resolution far beyond phenomenology. Other works refrain from modeling altogether[162, 163, 171].
What drew Bloembergen to this phenomenon — culminating in his SHCR trilogy[159, 165, 166] — and yet has defeated all challengers since, himself included? SHCR is nothing less than a Holy-Grail level[174] modeling challenge as an open benchmark in NCO computation, demanding equally deep command of LCO and NCO alike. The absence of either renders the process inherently intractable.
@@ 2.2 The union of CO, NO, and FO: a theoretical, experimental, and computational Holy Grail
On a phenomenological level, this NCO process(i.e. SHCR) involves synchronous but distinct LCO anisotropic diffraction for both the fundamental and harmonic waves(HWs). In Fig. 1f, the FW() conical refracts(CR) along its optic axis, while the HW() tends to double refract(DR) (yet mixed with CR, see Fig. S9 in Supplementary Note 8), forming a CRωCR-DR[134])2ω combination. In Fig. 2, pumped along the harmonic’s axis, the configuration shifts to DRωDR-CR[175])2ω. Each Fourier component of each eigenmode, in both 2D spatial and one-dimensional(1D) angular frequency domains , exhibits wave vector double(DR)/conical(CR)/double-conical(DR-CR) or even double conical refraction(DCR) (see Fig. 8d) at the material interfaces, superimposed on walk-off between Poynting vector and wave vector inside material.
Challenge deepens: what if the material is also chiral[144, 163, 162, 176, 160, 133, 177] and dichroic[134, 21, 22, 160, 139, 178, 157, 81], beyond being birefringent[22, 179, 134]? The nonlinear optics(NO) community has 0 solution for this — and rightly so. It lies at the pinnacle of linear crystal optics(LCO), a domain they are largely unfamiliar with, shaped over nearly 200 years by the world’s leading theorists[94, 180, 92, 181, 182, 183, 181, 184, 75, 185, 186, 187, 188, 189, 190, 191, 192, 5, 20, 101, 193, 194, 195], experimentalists[158, 165, 159, 196, 133, 197, 198, 199, 200, 201, 181, 202, 139, 178, 157, 81, 203, 204], applied[205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229], and computational[230, 95, 231, 46, 94, 96, 98, 232, 90, 211, 44, 233, 234, 235, 236, 237, 238, 54, 239, 240, 241, 62, 89, 91, 242, 243, 244] physicists.
@@
So, we ask, sincerely and directly, to the whole NO community, how, exactly, do you propose to calculate the Chiral SHCR process in Fig. 1? We assert, clearly, that the NO community currently have no means, by all means, none whatsoever, precisely because they have ignored the foundation of NCO — LCO, and unceasingly phenomenologically to decouple LCO(anisotropic diffraction) from NCO(frequency conversion processes), both of which occur in parallel across all spatiotemporal spectral components, all eigenmodes, and all tensor & vector components. Some ask, parallel computation? Yes, but using super-computers to solve tiny-problems?
What, then, is the true root cause? The above concerns only appearances. One phenomenon admits many explanations, a class of phenomena fewer, and all phenomena nearly one. That mechanism, though its initial discovery may be metaphysical, its ultimate core — mathematics — is anything but.
On a computational level, each of the six NCO phase-matching types() in Fig. 1a,d (and Fig. 2b) is, for chiral SHCR, phase-mismatched and “very likely” non-degenerate. Wherein, phase mismatch implies that none of these processes dominates the NCO in terms of energy conversion efficiency, necessitating the simultaneous consideration of all six. Non-degeneracy means that outcomes of the six (eigen)mode combinations() are usually distinct(see Fig. 1d), requiring separate calculations for each.
@@
For KTP, one vector NFCO process takes scalar NFCO runs (Fig. 1c). If one further scans wavelength (pulse injection, see Figs. 3, 1, 2 and 4 and SC2), propagation distance (dynamical evolution, see Figs. S9-S11, Fig. 8 and Fig. 1c), pump power, beam waist (see Fig. S11c), incident angle (see Fig. 2, Fig. 4, and Fig. S9), N.A. (see Fig. 8 and Fig. S11c), temperature , external magnetic field , -frame orientation (see Figs. SA4-SA5), or material coefficients (for adiabatic tuning, see Fig. 7 and Fig. S8) — each adds a for-loop layer, and each layer may grow due to low crystallographic symmetry, or dense parameter sampling.
@@
Therefore, to comprehensively investigate the high-dimensional parameter space of light-matter interactions, as done by Berry et al.(32 pages[5], 28 pages[20]), Mcleod et al.(45 pages[91]), A.Favaro et al.(54 pages[187]), Hehl et al.(25 pages[258]), this study, and many others[184, 259, 260, 261, 195, 174, 152], both linear and nonlinear crystal optics(L/NCO) call for theories/models/algorithms with a high speed-accuracy product.
@@
The incoming second part of this trilogy (our scalar NFCO model) points out: with a : scale mismatch — crystal macro(10 mm[162]), lightwave meso(1 m) — Green-function formalism[262, 263] struggles transversely/in-plane(x-y), while split-step Fourier[264, 265, 266, 267, 268, 269] and pseudo-spectral method[270, 271, 272] (with Runge-Kutta scheme[273, 44, 274, 275]) strain longitudinally/out-of-plane(z), making forward computation of even one scalar NFCO hard, let alone all 42 in a full-vector NFCO. To thoroughly resolve the scalar NFCO process, we formulate the imminent Nonlinear Angular Spectrum Theory(NLAST), with its validity demonstrated in Fig. 4 (and Figs. SB2-SB6).
@@
All forms of N(F)CO modeling — scalar (Fig. 4 and Figs. SB2-SB6), semi-vector (Fig. 5 and Figs. SC2), or full-vector (Figs. 1, 2 and 3 and Figs. SC3-SC5) — depend on a vector L(F)CO model (Figs. 6, 7 and 8, Figs. S8-S11, and Figs. SA2-SA10) to yield the requisite eigenvectors (eigen-polarizations) and eigenvalues (eigen-wavevector-z-components for FO) , necessary for evaluating the initial complex amplitudes (i.e. modal coefficients) in Eq. (S56c) to calculate frequency-mixing dynamics, and for determining the nonlinear conversion efficiency “eigenvalue mask” sinc “eigenvector mask” in Fig. 5.
@@
To showcase the basics of semi-vector NFCO, Fig. 5j reproduces the full conical phase matching(FCPM) second harmonic generation(SHG) proposed by Belyi et al.[221], illustrating the calculation of conversion in a uniaxial BBO crystal through Fig. 5a-i, where both the first- and second-order susceptibilities are anisotropic.
@@
In subsequent (both scalar and vector) NFCO, transverse wave vector conservation (Fig. 5a,b) is elevated to the same fundamental level as conservation and phase continuity at the boundaries, making it a prerequisite that must take priority. The conversion efficiency, determined solely by the longitudinal phase mismatch (Fig. 5f) and effective nonlinear coefficient (Fig. 5i)[277, 13, 3, 4]
@@
is subordinated to the fulfillment of transverse momentum conservation. Notably, , , and in Eq. 1 represents the eigen-polarization states of , , and in the frame, whose combination are fixed under certain phase-matching type. For example, phase-matching type-I ooe from Fig. 5g,h needs to compute = from Eq. 1. Above -frame eigenvectors are almost directly provided by Berry & Dennis’s 2003 model[5] in Fig. 10b, though a few minor adjustments in Fig. 10c are still required.
@@
As two masks determining the angular distribution of frequency conversion efficiency, the degree of longitudinal phase matching sinc from Fig. 5c and the effective nonlinear coefficient from Fig. 5i, respectively, depend on the eigenvalues and eigenvectors in linear Fourier crystal optics(LFCO), reaffirming that NCO is fundamentally grounded in LCO, and both linear(LO) and nonlinear optics(CO) should ultimately be incorporated into the framework of Fourier optics(FO).

### v6

@@ Abstract
To find a suitable framework for nonlinear crystal optics(NCO), we have revisited linear crystal optics(LCO). At the methodological level, three widely used plane wave bases are compared in terms of eigenanalysis in reciprocal space and light field propagation in real space. Inspired by complex ray tracing, we expand M.V. Berry & M.R. Dennis’s 2003 uniform plane wave model to non-uniform Fourier crystal optics(FCO) and ultimately derive the explicit form of its 32 transition matrix, bridging the two major branches of crystal optics(CO) in reciprocal space, where either ray direction or spatial frequency serves as the input variable. Using this model, we create the material-matrix tetrahedral compass(M-M TC) to conduct a detailed analysis of how the four fundamental characteristics of materials (linear/circular birefringence/dichroism) influence the eigensystems of the vector electric field in two-dimensional(2D) spatial frequency domain and its distribution in three-dimensional(3D) space with a crystal-2f configuration. Along this journey, we have uncovered new territories in LCO in both real and reciprocal space, such as infinite singularities arranged in disk-, ring-, and crescent-like shapes, “L shorelines” resembling hearts, generalized haunting theorem, double conical refraction(DCR), and optical knots(OKs) it induces. We also present our model’s early applications in focal engineering and NCO. For focusing: dual-eigenmode (1) decomposition, explaining Raman spikes in conical refraction(CR); (2) interference, leading to longitudinal multifoci; (3) backward propagation, for non-Zernike aberration correction. For NCO: (1) chiral second harmonic conical refraction(SHCR); (2) linear and nonlinear spin-orbit coupling; (3) nonlinear optical(NO) orbital angular momentum(OAM) non-conservation — all three vector/scalar NCO phenomena, reproduced pixel by pixel. As the opening chapter in a trilogy, this work interweaves crystal optics, Fourier optics, and nonlinear optics, while integrating theoretical, computational, and experimental physics, advancing all six domains.
@@
Crystal optics, Singular optics, Fourier optics, Nonlinear optics, Quantum opticsContents
@@ 1 Introduction
Threefold purpose guides us: to establish crystal optics(CO) and Fourier optics(FO) as the physical and computational bases of nonlinear crystal optics(NCO) in the first part of a trilogy; to propose that three plane-wave bases form the mathematical core of CO spectral analysis, each offering trade-offs; to seek and knock the computational boundaries of linear crystal optics(LCO) and NCO via large-scale FO simulations in both three-dimensional(3D) reciprocal and two-dimensional(2D) spatial domains.
@@
Think about the ordinary(o) and exordinary(e) waves in nonlinear crystal optical(NCO) phase-matching types[6, 7, 8, 9, 10, 11, 12](e.g. type-I ooe in Fig.˜5a): what are they? Eigenmodes governed by CO[13, 4, 14, 15, 16, 17, 18, 19]!! That is to say, the entire NCO process is constantly subject to the constrains imposed by CO[13, 5, 3, 4], which is itself inherently complex[20]. More phenomenologically, for all frequency-mixing phenomena, both the pumps/fundamental waves(FWs) and the newly generated frequencies(NFs) within the crystal, must be first decomposed into the material’s intrinsic eigenmodes, then diffract independently (and anisotropically) during subsequently linearly superposition and interference, while potentially enduring absorption or gain[21, 22].
@@
Given that the impact of LCO(= LO CO) on NCO has been historically either ignored or modeled incorrectly[22], we undertake a thorough reexamination of LCO in Supplementary Note 1. Through this inquiry, we summarize that:
@@
Only two options remain (see “Methods”): uniform CO basis that contradicts both FO and boundary conditions(BCs), but only requires solving bi-quadratic equations thanks to spherical coordinates()[5]; rectangular() non-uniform CO basis that fits both physics(FO) and experiments(BCs), but demands batchly solving quartics with intractably long formulas[96].
@@
In this Article, we propose a non-uniform linear Fourier crystal optical(LFCO) model, whose CO part integrates the advantages of both uniform[75, 76, 77, 78, 79, 80, 81, 1, 5, 20] and non-uniform[83, 15, 85, 14, 84, 86, 87, 111, 112] plane wave models , solving a pure bi-quadratic under FO and boundary constraints. This LFCO model extends the standard non-diagonal/non-positive-definite[113, 114, 115, 116, 117, 118, 119, 113, 120, 121, 122]/non-Hermitian[123, 124, 125, 126, 127, 128, 129, 130, 131, 132]/non-unitary[133, 134, 135, 136]/non-normal[107]/non-(mirror-[137])symmetric[5, 138, 139]/non-reciprocal[140, 141, 137] 22[142] transfer[98, 143]/transition[98, 110] matrix( evolution operator[144]) in CO to a 32 form (see Eqs.˜5 and 6), enabling explicit fully vectorial[93, 91, 145, 146, 147, 148, 149, 150, 151] computation within arbitrary materials, and at the same time, allows the LO part to directly operate in the non-uniform[83, 15, 85, 14, 84, 86, 87, 92, 84] FO framework[91, 62, 90, 93, 89].
@@
To indirectly validate this LFCO model, and highlight its powerful role in advancing the analytic development of NCO, we offer a roadmap for expanding the LFCO model in Fig.˜2c. Following its conceptual architectures — scalar(, semi-vector,) and (full-)vector NFCO — we first demonstrate this LFCO model’s applications: in full-vector NFCO, via chiral second-harmonic conical refraction(SHCR) in Figs.˜1 and 2, and harmonic spin-orbit(S-O) angular momentum(AM) cascade in Fig.˜4 (see also Figs. SC3-SC5); in scalar NFCO, through phase-matching-controlled orbital angular momentum(OAM) conversion in Fig.˜3 (see also Figs. SB2-SB6); and in semi-vector NFCO, via full conical phase matching(FCPM) in Fig.˜5 (see also Fig. SC2).
@@
To directly validate this LFCO model, as stated at the beginning of Section˜2.4, the procedure must begin with verifying the correctness of the eigenvalue-eigenvector computations (Fig. S8), followed by confirming the accuracy of the field distributions in real space (Fig.˜7, Fig.˜8, Figs. S9-S11, and Figs. SA2-SA10).
@@
Accordingly, we introduce the material-matrix(M-M) tetrahedron compass(TC) in Fig. S8, where scanning parameters along three of its edges depict a theoretical panorama of the adiabatic evolution(AE) of electric field eigenmodes in the domain[152], competing among the three primary material properties, i.e. linear dichroism(LD), circular dichroism(CD), and optical activity(OA). During this process, we observe that CD, adhering to the haunting (C points) theorem[5] as well, can lead to heart-shaped L shorelines (as an upgrade to the L lines in CO[5]) and an infinite array of singularities in 2D domain, often arranged in patterns resembling disks, rings[101, 153], or crescents (instead of 8 finite EPs[5, 154, 154, 155, 22, 92, 156]). In Fig. S8, certain known phenomena reappeared[101, 153, 157], while the others, to the best of our knowledge, are originally predicted.
@@
Aided by crystal-2f system proposed from Fig. S9, we map out a second, more experimentally relevant[158, 159, 139, 81, 160] panorama in the 3D domain in Fig.˜7, freezing the AE of the light field distribution between the three vertices of the M-M TC, corresponding to the material’s birefringence(Bi), LD, and OA.
@@
Still in real space, but with higher numerical aperture(N.A.), we offer a unified solution for the forward propagation and inverse design of focal fields in Fig.˜8, where a dual-eigenmode decomposition is proposed to explain Raman spikes[161] in conical refraction(CR) arising solely from slow modes, multifoci induced by laser processing inside materials, and aberration correction without Zernike polynomials. Along the way, we have observed two surprising new phenomena: double conical refraction(DCR) and the resulting optical field knots.
@@
This work unites mathematical elegance(see Supplementary Notes 2-5), physical intuition(see Supplementary Notes 7), and experimental accuracy(see Supplementary Note 8-9). Even so, it too currently suffers computational limits(see Supplementary Notes 5-6).
@@ 2 Results
2.1 By 2025, it still remains a great challenge to reproduce chiral SHCR
@@
The optically active( chiral[160]) SHCR[159, 165, 166, 162, 163] experiment in Fig.˜1f[162] represents an NCO experiment that, in principle, cannot be well-reproduced without a mature LCO framework[163, 164, 167, 168, 166, 169, 170, 171, 162, 172, 173]. Almost all previous NCO modeling attempts[170, 168, 168, 166, 167] have invariably failed to capture the experimental phenomena at a pixel-level resolution far beyond phenomenology. Other works refrain from modeling altogether[162, 163, 171].
As the concluding chapter[166] of Bloembergen’s trilogy[159, 165, 166], SHCR is nothing less than a Holy-Grail level[174] modeling challenge as an open benchmark in NCO computation, demanding equally deep command of LCO and NCO alike. The absence of either renders the process inherently intractable.
@@ 2.2 The union of CO, NO, and FO: a theoretical, experimental, and computational Holy Grail
On a phenomenological level, this NCO process(i.e. SHCR) involves synchronous but distinct LCO anisotropic diffraction for both the fundamental and harmonic waves(HWs). In Fig.˜1f, the FW() conical refracts(CR) along its optic axis, while the HW() tends to double refract(DR) (yet mixed with CR, see Fig. S9 in Supplementary Note 8), forming a CRωCR-DR[134])2ω combination. In Fig.˜2, pumped along the harmonic’s axis, the configuration shifts to DRωDR-CR[175])2ω. Each Fourier component of each eigenmode, in both 2D spatial and one-dimensional(1D) angular frequency domains , exhibits wave vector double(DR)/conical(CR)/double-conical(DR-CR) or even double conical refraction(DCR) (see Fig.˜8d) at the material interfaces, superimposed on walk-off between Poynting vector and wave vector inside material.
Challenge deepens: what if the material is also chiral[144, 163, 162, 176, 160, 133, 177] and dichroic[134, 21, 22, 160, 139, 178, 157, 81], beyond being birefringent[22, 179, 134]? The corresponding LCO-based NCO process lies at the pinnacle of linear crystal optics(LCO), a domain shaped over nearly 200 years by the world’s leading LCO theorists[94, 180, 92, 181, 182, 183, 181, 184, 75, 185, 186, 187, 188, 189, 190, 191, 192, 5, 20, 101, 193, 194, 195], experimentalists[158, 165, 159, 196, 133, 197, 198, 199, 200, 201, 181, 202, 139, 178, 157, 81, 203, 204], applied[205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229], and computational[230, 95, 231, 46, 94, 96, 98, 232, 90, 211, 44, 233, 234, 235, 236, 237, 238, 54, 239, 240, 241, 62, 89, 91, 242, 243, 244] physicists.
@@
So, we ask, how, exactly, to calculate the Chiral SHCR process in Fig.˜1? Most previous efforts have ignored the foundation of NCO — LCO, and unceasingly phenomenologically to decouple LCO(anisotropic diffraction) from NCO(frequency conversion processes), both of which occur in parallel across all spatiotemporal spectral components, all eigenmodes, and all tensor & vector components. Some ask, parallel computation? Yes, but still either mathematically incorrect or numerically inefficient.
On a computational level, each of the six NCO phase-matching types() in Fig.˜1a,d (and Fig.˜2b) is, for chiral SHCR, phase-mismatched and “very likely” non-degenerate. Wherein, phase mismatch implies that none of these processes dominates the NCO in terms of energy conversion efficiency, necessitating the simultaneous consideration of all six. Non-degeneracy means that outcomes of the six (eigen)mode combinations() are usually distinct(see Fig.˜1d), requiring separate calculations for each.
@@
For KTP, one vector NFCO process takes scalar NFCO runs (Fig.˜1c). If one further scans wavelength (pulse injection, see Figs.˜4, 1, 2 and 3 and SC2), propagation distance (dynamical evolution, see Figs. S9-S11, Fig.˜8 and Fig.˜1c), pump power, beam waist (see Fig. S11c), incident angle (see Fig.˜2, Fig.˜3, and Fig. S9), N.A. (see Fig.˜8 and Fig. S11c), temperature , external magnetic field , -frame orientation (see Figs. SA4-SA5), or material coefficients (for adiabatic tuning, see Fig.˜7 and Fig. S8) — each adds a for-loop layer, and each layer may grow due to low crystallographic symmetry, or dense parameter sampling.
@@
Therefore, to comprehensively investigate the high-dimensional parameter space of light-matter interactions, as done by Berry et al.[5, 20], Mcleod et al.[91], A.Favaro et al.[187], Hehl et al.[258], this study, and many others[184, 259, 260, 261, 195, 174, 152], both linear and nonlinear crystal optics(L/NCO) call for theories/models/algorithms with a high speed-accuracy product.
@@
The incoming second part of this trilogy (our scalar NFCO model) points out: with a : scale mismatch — crystal macro(10 mm[162]), lightwave meso(1 m) — Green-function formalism[262, 263] struggles transversely/in-plane(x-y), while split-step Fourier[264, 265, 266, 267, 268, 269] and pseudo-spectral method[270, 271, 272] (with Runge-Kutta scheme[273, 44, 274, 275]) strain longitudinally/out-of-plane(z), making forward computation of even one scalar NFCO hard, let alone all 42 in a full-vector NFCO. To thoroughly resolve the scalar NFCO process, we formulate the imminent Nonlinear Angular Spectrum Theory(NLAST), with its validity demonstrated in Fig.˜3 (and Figs. SB2-SB6).
@@
All forms of N(F)CO modeling — scalar (Fig.˜3 and Figs. SB2-SB6), semi-vector (Fig.˜5 and Figs. SC2), or full-vector (Figs.˜1, 2 and 4 and Figs. SC3-SC5) — depend on a vector L(F)CO model (Figs.˜6, 7 and 8, Figs. S8-S11, and Figs. SA2-SA10) to yield the requisite eigenvectors (eigen-polarizations) and eigenvalues (eigen-wavevector-z-components for FO) , necessary for evaluating the initial complex amplitudes (i.e. modal coefficients) in Eq. (S56c) to calculate frequency-mixing dynamics, and for determining the nonlinear conversion efficiency “eigenvalue mask” sinc “eigenvector mask” in Fig.˜5.
@@
To showcase the basics of semi-vector NFCO, Fig.˜5j reproduces the full conical phase matching(FCPM) second harmonic generation(SHG) proposed by Belyi et al.[221], illustrating the calculation of conversion in a uniaxial BBO crystal through Fig.˜5a-i, where both the first- and second-order susceptibilities are anisotropic.
@@
In subsequent (both scalar and vector) NFCO, transverse wave vector conservation (Fig.˜5a,b) is elevated to the same fundamental level as conservation and phase continuity at the boundaries, making it a prerequisite that must take priority. The conversion efficiency, determined solely by the longitudinal phase mismatch (Fig.˜5f) and effective nonlinear coefficient (Fig.˜5i)[277, 13, 3, 4]
@@
is subordinated to the fulfillment of transverse momentum conservation. Notably, , , and in Eq.˜1 represents the eigen-polarization states of , , and in the frame, whose combination are fixed under certain phase-matching type. For example, phase-matching type-I ooe from Fig.˜5g,h needs to compute = from Eq.˜1. Above -frame eigenvectors are almost directly provided by Berry & Dennis’s 2003 model[5] in Fig.˜10b, though a few minor adjustments in Fig.˜10c are still required.
@@
As two masks determining the angular distribution of frequency conversion efficiency, the degree of longitudinal phase matching sinc from Fig.˜5c and the effective nonlinear coefficient from Fig.˜5i, respectively, depend on the eigenvalues and eigenvectors in linear Fourier crystal optics(LFCO), reaffirming that NCO is fundamentally grounded in LCO, and both linear(LO) and nonlinear optics(NO) should ultimately be incorporated into the framework of Fourier optics(FO).